# DanModCap: Designing a Danmaku Moderation Tool for Video-Sharing Platforms that Leverages Impact Captions

### v2

@@ Abstract.
Figure 1. Our Danmaku moderation approach. 1) Typical Danmaku involves positive, inappropriate, or irrelevant content. 2) On video-sharing platforms, Danmaku is a built-in social feature that supports end-users’ social communication and is synchronized with the video. 3) To moderate Danmaku content, we introduce a proactive approach called Impact Captions, that leverages a key feature in East Asian variety shows. 4) Building on this conceptual foundation, we introduce DanModCap, a tool for Danmaku content moderation that uses generative AI to create Impact Captions. 5 and 6) The pipeline of Impact Caption generation and example captions, which include the Tasukkomi style (red), the Explanatory style (blue), and the Humorous and Praising style (orange).
@@ 1. Introduction
This research first conducted a video-content analysis to understand the existing uses of Impact Caption design. We identified a taxonomy of Impact Caption types that was organized into three categories and seven design dimensions. We then ran an expert-based co-design workshop to solicit design recommendations to steer Danmaku content toward the promotion of prosocial behaviors. Spurred by the expert recommendations and our taxonomy, we developed a Danmaku moderation tool, DanModCap, that would support users in moderating live content while being mindful of community engagement and the dynamic nature of viewer pro-social interactions. A key tenant of DanModCap is the Impact Caption, a popular feature in East Asian variety shows (Basaran and Dilber, 2013; Sasamoto, 2014). Impact Captions are visually compelling and context-sensitive textual and non-textual elem that are designed to captivate the reader’s attention and enhance their emotional experience (Achlioptas et al., 2021). They are powerful triggers that add depth and meaning to content. Our implementation of Impact Captions was rooted in the resonance theory  (Decety and Meyer, 2008; McDonnell et al., 2017), which suggests that when we see or hear something that resonates with us emotionally and cognitively, it creates a stronger connection to us. By using familiar phrases to encourage cognitive resonance (Kirsh, 2013) and vivid colors and speech bubbles to encourage emotional resonance (Wan et al., 2024), Impact Captions have the potential to bridge the gap between users and content, thus helping to shape more prosocial behaviors on VSPs.
@@
To explore the influence of the generative Impact Captions intervention on viewers’ internal emotions, cognition, attitudes, and external social behaviors, we then conducted a user study. The study found that viewers experienced both cognitive and emotional resonance with the generative Impact Captions, perceiving them as aligned with their own thoughts, experiences, and intentions – particularly around online activism and community belonging. The captions influenced viewers’ attention dynamics, at times diverting focus from the main video content to the captions themselves. They also fostered a heightened sense of community cohesion, belongingness, companionship, and security, among other findings.
Due to these research activities, we thus contribute:
@@ 2.2. Content Moderation Approaches
Algorithms like Latent Dirichlet Allocation (LDA) have also been shown to provide more accurate and efficient content moderation (Vörös et al., 2023; Kolla et al., 2024). LDA is often used to extract hidden topic information from large amounts of text data as a topic modeling technology. Ammari et al., for example, transformed Reddit threads into a Bag of Words model and applyied LDA to identify distinct topics that parents were likely to discuss anonymously (Ammari et al., 2019) . Jhaver et al. utilized LDA to analyze the content of removal explanations on Reddit, aiming to understand how these communications impact user behavior (Jhaver et al., 2019). Inspired by this research, we used LDA to extract hidden thematic information from Danmaku comment text to identify the core themes of user discussions.
@@ 3. Study 1: Impact Caption Video Analysis
3.1. Procedure
We initially consulted three global internet film and television databases, i.e., IMDB [^1] IMDB: https://www.imdb.com/, TV Time [^2] TV Time: https://www.tvtime.com/, and Douban [^3] Douban: https://www.douban.com/, to identify highly popular variety shows from East Asia that had an airing history of at least eight years and had consistently incorporated Impact Caption content. These included China’s Who is the Murderer [^4] Who is the Murderer: https://en.wikipedia.org/wiki/Who%27s_the_Murderer, South Korea’s Running Man [^5] Running Man: https://en.wikipedia.org/wiki/Running_Man_(TV_video), and Japan’s VS ARASHI [^6] VS ARASHI: https://zh.wikipedia.org/wiki/VS%E5%B5%90. We then watched the top three episodes from each show that had the highest viewership to gain a preliminary understanding of the show’s content. While watching, we identified that Impact Captions had several diverse visual design features that influenced audience appeal and manipulation based on previously work (Sasamoto, 2014). For example, we found that videos with Impact Caption employed a series of striking visual design elements, including vivid color contrasts, dynamic typography layouts, and theme-coordinated background images.
Following this, we examined YouTube [^7] YouTube: https://www.youtube.com/ and Bilibili [^8] Bilibili: https://www.bilibili.com/ to analyze current design trends and the utilization of Impact Captions in variety shows on these platforms as online video platforms allow for more experimentation and innovation in visual design compared to more constrained television formats. Three researchers conducted keyword searches on these video-sharing platforms (e.g., “variety show Impact Caption”, “Impact Caption in the TV program” and “video Impact Caption”) to find and select representative samples based on the number of views. After filtering out irrelevant and duplicate content, 27 videos with an average duration of 43.23 minutes and 33.83k views were selected for analysis. The researchers then conducted independent open coding of a subset of the videos to identify critical aspects of Impact Captions, including visual elements, interaction patterns, and the perspective, using methods established in prior work ( (Sasamoto, 2014; Lu and Lu, 2022)). Following multiple rounds of reflection and discussion, a taxonomy of Impact Captions was developed, with three types of captions that spanned seven design dimensions (Figure 2). The outcomes from this analysis were used to inform our subsequent research activities.
@@ 3.2.2. Caption Style 2: Multi-Perspective Information
With Multi-Perspective Information, two perspectives were used: first person points of view (POV) and third person POV. First person POVs centered on the protagonist’s perspective, directly disclosing their experiences, thoughts, actions, and emotions. These captions were instrumental in forging an empathetic connection between the lead character and the audience. In contrast, third person POVs offered a more comprehensive, ‘removed’ viewpoint, encompassing additional details, commentaries, and humorous response styles like Tsukkomi (Roasting), Expository or Humorous and Praise. This approach significantly broadened the entertainment value and deepened audience understanding, offering an enriched viewing experience.
@@ 4.1. Participants
We recruited six participants for our co-design workshop. The group was comprised of three males and three females (all aged between 27 and 35) with backgrounds spanning broadcasting, television editing and directing, and visual design Participants included motion graphic designers, video editors, and post-production supervisors, each of whom had hands-on experience incorporating video impact captions into their work. To avoid bias, our recruitment process intentionally omitted terms like “variety show Impact Caption” and “Danmaku”, instead framing the work as “Design and Research on Video Sharing Platforms”. While the small number of participants could be considered a potential limitation, the depth of their industry expertise and specialized knowledge provided rich, context-specific insights that informed the subsequent stages of our research.
@@ 5.1. Design Goals
  • G1: Support the proactive moderation of Danmaku through textual and non-textual Impact Caption elements to improve the efficiency of curation and reduce the need for post-video processing.
  • G2: Encourage spontaneous prosocial user behavior for Danmaku-streaming content to reduce moderation burdens and improve the overall user experience.
  • G3: Construct cognitive and emotional resonance via textual and non-textual elements while being mindful of relevance, cultural sensitivity, and visual appeal.
With G1, we sought to actively manage Danmaku content on video platforms via several methods. First, by implementing an AI-based automatic content filtering system we wanted to analyze the content and emotional tendencies of the Danmaku in real time and identify and filter out undesired words, offensive speech, hateful speech, and inappropriate jokes. Second, the creation of a real-time monitoring mechanism, especially during peak hours or when popular programs are broadcast, would allow us to quickly identify and deal with problematic Danmaku information. Lastly, we wanted to understand community norms so we could convey clear Danmaku usage norms.
With G2, we wanted to guide users to actively abide by community norms by promoting positive and responsible social behaviors. We sought to realize this in several ways. First, through educational activities and guidance, we wanted to help users realize and understand the importance of community norms, encourage them to abide by these norms, and educate them on how to use Danmaku with responsibility and respect. Reward mechanisms (e.g., positive feedback or responses such as echoes and praises) were also designed to recognize users and comments that show positive prosocial behaviors. Users can be encouraged to participate in the process of Danmaku supervision by bringing personal perspectives to the role of community moderator and encouraging positive interactions between users to reduce negative or harmful Danmaku content.
With G3, we will harness Impact Captions to nurture prosocial behavior within video platform communities. Our captions will be crafted to evoke emotional resonance through relatable text and graphics to encourage active community engagement, offering a subtle nudge towards thoughtful participation. Celebrating positive contributions through these captions will not only highlight model behavior but also set a communal standard, motivating users to contribute to a respectful and engaging dialogue.
@@ 5.2. Tenants of Impact Captions
UTF8gbsn Tsukkomi (Roasts) style is utilized with negatively oriented Danmaku content. It leverages an informal communication approach, commonly identifying loopholes or keywords in another’s language or behavior, followed by the expression of irony or doubt. According to interpersonal communication theory, this style employs humor to address critical comments while avoiding excessive negativity, thus preventing the treatment of viewers as “emotional garbage” (Hofmann et al., 2016). For instance, hyperbolic metaphors might be used to humorously address narcissistic tendencies, like “这位自恋狂是不是以为自己是全宇宙的中心啊?你有皇位要继承吗? (Does this narcissist think they are the center of the universe? Do they have a throne to inherit?)”. This style facilitates the collective criticism of disruptive Danmaku content, promoting emotional regulation through shared complaints, which can engender a sense of unity and intimacy among strangers (Suls et al., 2002; Ellemers et al., 2002).
@@ 5.3.1. Step 1: Identifying Danmaku Topics and Sentiment
To identify Danmaku behavior, the content and context of user discussions must be captured within a designated time interval and then the sentiment and expressed emotions must be evaluated to discern their behavioral patterns. Utilizing Latent Dirichlet Allocation (LDA)  (Blei et al., 2003), the main topics from each Danmaku were extracted. LDA was chosen due its suitability to handle textual data, which is necessary for Danmaku comments. This method is particularly apt for scenarios involving large volumes of open-ended comment data, as it allows for the identification of the topics within the data without the need for predefined categories of themes. Once the topics were identified, a multi-lingual fine-tuned BERT model  (Pires et al., 2019), trained on the GoEmotions dataset with 28 sentiment labels, was used to categorize the emotions expressed in the topics. This approach facilitated the comprehension of the content and sentiment of Danmaku discussions.
@@ 5.3.3. Step 3: Generating Impact Caption Text and Graphics
By integrating LDA, LLaMA2, and Stable Diffusion, DanModCap could understand and modulate the content of Impact Captions to handle the rapid flow Danmaku and its cultural nuances. By using data pre-processing, topic extraction, and sentiment analysis (Step 1), alongside text generation (Step 3), and the intelligent design of visual elements (Step 3), DanModCap should enable Impact Captions to resonate emotionally and visually with the audience.
@@ 7.3. Risks with Automated Moderation Methods
Finally, an over-reliance on AI generated content poses a significant risk in relation to consistency and accuracy output (Chen and Shu, 2023; Uchendu et al., 2023). The potential inaccuracies that can emerge while trying to generate complex human emotions and cultural subtleties, coupled with the reduced level of human creativity and authorship, can lead to undesirable user experiences.

### v3

@@ Abstract.
Figure 1. Our Danmaku moderation approach. 1) Typical Danmaku involves positive, inappropriate, or irrelevant content. 2) On video-sharing platforms, Danmaku is a built-in social feature that supports end-users’ social communication and is synchronized with the video. 3) To moderate Danmaku content, we introduce a proactive approach called Impact Captions, that leverages a key feature in East Asian variety shows. 4) Building on this conceptual foundation, we introduce DanModCap, a tool for Danmaku content moderation that uses generative AI to create Impact Captions. 5 and 6) The pipeline of Impact Caption generation and example captions, which include the Tsukkomi style (red), the Explanatory style (blue), and the Humorous and Praising style (orange).
@@ 1. Introduction
This research first conducted a video-content analysis to understand the existing uses of Impact Caption design. We identified a taxonomy of Impact Caption types that was organized into three categories and seven design dimensions. We then ran an expert-based co-design workshop to solicit design recommendations to steer Danmaku content toward the promotion of prosocial behaviors. Spurred by the expert recommendations and our taxonomy, we developed a Danmaku moderation tool, DanModCap, that would support users in moderating live content while being mindful of community engagement and the dynamic nature of viewer pro-social interactions. A key tenant of DanModCap is the Impact Caption, a popular feature in East Asian variety shows (Basaran and Dilber, 2013; Sasamoto, 2014). Impact Captions are visually compelling and context-sensitive textual and non-textual elements that are designed to captivate the reader’s attention and enhance their emotional experience (Achlioptas et al., 2021). They are powerful triggers that add depth and meaning to content. Our implementation of Impact Captions was rooted in the resonance theory  (Decety and Meyer, 2008; McDonnell et al., 2017), which suggests that when we see or hear something that resonates with us emotionally and cognitively, it creates a stronger connection to us. By using familiar phrases to encourage cognitive resonance (Kirsh, 2013) and vivid colors and speech bubbles to encourage emotional resonance (Wan et al., 2024), Impact Captions have the potential to bridge the gap between users and content, thus helping to shape more prosocial behaviors on VSPs (Video-Sharing Platforms).
@@
To explore the influence of the generative Impact Captions intervention on viewers’ internal emotions, cognition, attitudes, and external social behaviors, we then conducted a user study. The study found that viewers experienced both cognitive and emotional resonance with the generative Impact Captions, perceiving them as aligned with their own thoughts, experiences, and intentions – particularly around online activism and community belonging. The captions influenced viewers’ attention dynamics, at times diverting focus from the main video content to the captions themselves. They also fostered a heightened sense of community cohesion, belongingness, companionship, and security, among other findings. Due to these research activities, we thus contribute:
@@ 2.2. Content Moderation Approaches
Algorithms like Latent Dirichlet Allocation (LDA) have also been shown to provide more accurate and efficient content moderation (Vörös et al., 2023; Kolla et al., 2024). LDA is often used to extract hidden topic information from large amounts of text data as a topic modeling technology. Ammari et al., for example, transformed Reddit threads into a Bag of Words model and applied LDA to identify distinct topics that parents were likely to discuss anonymously (Ammari et al., 2019) . Jhaver et al. utilized LDA to analyze the content of removal explanations on Reddit, aiming to understand how these communications impact user behavior (Jhaver et al., 2019). Inspired by this research, we used LDA to extract hidden thematic information from Danmaku comment text to identify the core themes of user discussions.
@@ 3.1.2. Caption Style 2: Multi-Perspective Information
With Multi-Perspective Information, two perspectives were used: first person point of view (POV) and third person POV. First person POVs centered on the protagonist’s perspective, directly disclosing their experiences, thoughts, actions, and emotions. These captions were instrumental in forging an empathetic connection between the lead character and the audience. In contrast, third person POVs offered a more comprehensive, ‘removed’ viewpoint, encompassing additional details, commentaries, and humorous response styles like Tsukkomi (Roasting), Expository or Humorous and Praise. This approach significantly broadened the entertainment value and deepened audience understanding, offering an enriched viewing experience.
@@ 4.1. Participants
We recruited six participants for our co-design workshop. The group was comprised of three males and three females (all aged between 27 and 35) with backgrounds spanning broadcasting, television editing and directing, and visual design. Participants included motion graphic designers, video editors, and post-production supervisors, each of whom had hands-on experience incorporating video impact captions into their work. To avoid bias, our recruitment process intentionally omitted terms like “variety show Impact Caption” and “Danmaku”, instead framing the work as “Design and Research on Video Sharing Platforms”. While the small number of participants could be considered a potential limitation, the depth of their industry expertise and specialized knowledge provided rich, context-specific insights that informed the subsequent stages of our research.
@@ 5.1. Design Goals
  • DG1: Support the proactive moderation of Danmaku through textual and non-textual Impact Caption elements to improve the efficiency of curation and reduce the need for post-video processing.
  • DG2: Encourage spontaneous prosocial user behavior for Danmaku-streaming content to reduce moderation burdens and improve the overall user experience.
  • DG3: Construct cognitive and emotional resonance via textual and non-textual elements while being mindful of relevance, cultural sensitivity, and visual appeal.
With [DG1], we sought to actively manage Danmaku content on video platforms via several methods. First, by implementing an AI-based automatic content filtering system we wanted to analyze the content and emotional tendencies of the Danmaku in real time and identify and filter out undesired words, offensive speech, hateful speech, and inappropriate jokes. Second, the creation of a real-time monitoring mechanism, especially during peak hours or when popular programs are broadcast, would allow us to quickly identify and deal with problematic Danmaku information. Lastly, we wanted to understand community norms so we could convey clear Danmaku usage norms.
With [DG2], we wanted to guide users to actively abide by community norms by promoting positive and responsible social behaviors. We sought to realize this in several ways. First, through educational activities and guidance, we wanted to help users realize and understand the importance of community norms, encourage them to abide by these norms, and educate them on how to use Danmaku with responsibility and respect. Reward mechanisms (e.g., positive feedback or responses such as echoes and praises) were also designed to recognize users and comments that show positive prosocial behaviors. Users can be encouraged to participate in the process of Danmaku supervision by bringing personal perspectives to the role of community moderator and encouraging positive interactions between users to reduce negative or harmful Danmaku content.
With [DG3], we will harness Impact Captions to nurture prosocial behavior within video platform communities. Our captions will be crafted to evoke emotional resonance through relatable text and graphics to encourage active community engagement, offering a subtle nudge towards thoughtful participation. Celebrating positive contributions through these captions will not only highlight model behavior but also set a communal standard, motivating users to contribute to a respectful and engaging dialogue.
@@ 5.2. Tenants of Impact Captions
UTF8gbsn The Tsukkomi (Roasts) style is utilized with negatively oriented Danmaku content. It leverages an informal communication approach, commonly identifying loopholes or keywords in another’s language or behavior, followed by the expression of irony or doubt. According to interpersonal communication theory, this style employs humor to address critical comments while avoiding excessive negativity, thus preventing the treatment of viewers as “emotional garbage” (Hofmann et al., 2016). For instance, hyperbolic metaphors might be used to humorously address narcissistic tendencies, like “这位自恋狂是不是以为自己是全宇宙的中心啊?你有皇位要继承吗? (Does this narcissist think they are the center of the universe? Do they have a throne to inherit?)”. This style facilitates the collective criticism of disruptive Danmaku content, promoting emotional regulation through shared complaints, which can engender a sense of unity and intimacy among strangers (Suls et al., 2002; Ellemers et al., 2002).
@@ 5.3.1. Step 1: Identifying Danmaku Topics and Sentiment
To identify Danmaku behavior, the content and context of user discussions must be captured within a designated time interval and then the sentiment and expressed emotions must be evaluated to discern their behavioral patterns. Utilizing Latent Dirichlet Allocation (LDA)  (Blei et al., 2003), the main topics from each Danmaku were extracted. LDA was chosen due to its suitability to handle textual data, which is necessary for Danmaku comments. This method is particularly apt for scenarios involving large volumes of open-ended comment data, as it allows for the identification of the topics within the data without the need for predefined categories of themes. Once the topics were identified, a multi-lingual fine-tuned BERT model  (Pires et al., 2019), trained on the GoEmotions dataset with 28 sentiment labels, was used to categorize the emotions expressed in the topics. This approach facilitated the comprehension of the content and sentiment of Danmaku discussions.
@@ 5.3.3. Step 3: Generating Impact Caption Text and Graphics
By integrating LDA, LLaMA2, and Stable Diffusion, DanModCap could understand and modulate the content of Impact Captions to handle the rapid flow of Danmaku and its cultural nuances. By using data pre-processing, topic extraction, and sentiment analysis (Step 1), alongside text generation (Step 3), and the intelligent design of visual elements (Step 3), DanModCap should enable Impact Captions to resonate emotionally and visually with the audience.
@@ 7.3. Risks with Automated Moderation Methods
Finally, an over-reliance on AI generated content poses a significant risk in relation to consistency and accuracy of output (Chen and Shu, 2023; Uchendu et al., 2023). The potential inaccuracies that can emerge while trying to generate complex human emotions and cultural subtleties, coupled with the reduced level of human creativity and authorship, can lead to undesirable user experiences.